OpenStudy (openstudy9):

What is the Period of the function y=cos (x/5)? I know that the formula for period is 2pi/Absolute value of b.

OpenStudy (irishboy123):

here's a way to remember this if you like the normal formulation [in physics\engineering] is \(y = \cos \omega x\). \(\omega\) is the angular speed [or angular frequency] measured in radians per second using a simple distance-speed-time formula for angular motion, you can say that \(\omega = \dfrac{2 \pi}{T}\) where T is the period [for completeness, you can also say that \(\omega = 2 \pi f\) where fi is the ordinary frequency measured in Hertz ie oscillations per second] so here you have \(\omega = \dfrac{1}{5}\) and you can now do the rest.....

OpenStudy (openstudy9):

so 2pi/1.5 !?

OpenStudy (irishboy123):

from \(\omega = \dfrac{2 \pi}{T}\) you can say \(T = \dfrac{2 \pi}{\omega} = \dfrac{2 \pi}{\frac{1}{5}} = \dots\)